# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v1

@@ 3 Admissible Strategies
Note $V_{win}$, $V_{los}$, and $V_{pen}$ define a partition for $V$, i.e., their union is $V$ and their pair-wise intersection is the empty set. Based on these sets, we characterize value-preserving strategies according to the notion of value for each state. Let $\operatorname{sVal}:V\to\{-1,0,1\}$ be a state-value function such that $\operatorname{sVal}(v)=1$ if $v\in V_{win}$, 00 if $v\in V_{pen}$, $-1$ if $v\in V_{los}$.
@@
We say history $h$ is value-preserving if $\operatorname{sVal}(h_{j})\leq\operatorname{sVal}(h_{j+1})$ for all $10\leq j<|h|-$ 1. Strategy $\sigma$ is value preserving, if every $h\in\operatorname{Plays}^{v_{0}}(\sigma)$ is value preserving.
Let us now look at two classical notions for strategies defined for quantitative games and discuss their value-preserving property. We say $\sigma$ is a worst-case optimal strategy ($\operatorname{\mathbf{WCO}}$) if $\operatorname{aVal}(h,\sigma)=\operatorname{aVal}(h)$. If at the current state $v$, $\operatorname{sVal}(v)=1$, then an optimal winning strategy $\sigma_{win}$ exists such that all plays in $\operatorname{Plays}^{v}(\sigma_{win})$ are value-preserving. Since $\sigma_{win}$ is $\operatorname{\mathbf{WCO}}$, every $\operatorname{\mathbf{WCO}}$ strategy in $V_{win}$ is also value preserving. If $\operatorname{sVal}(v)\neq-1$, a cooperatively-optimal ($\operatorname{\mathbf{Co-Op}}$) strategy $\sigma$ exists such that $\operatorname{cVal}(h,\sigma)=\operatorname{cVal}(h)$. Unlike worst-case optimal strategies, $\operatorname{\mathbf{Co-Op}}$ are not value-preserving. In Fig. 2, $\sigma_{1}$ and $\sigma_{2}$ are $\operatorname{\mathbf{WCO}}$ while $\sigma_{3}$ is $\operatorname{\mathbf{Co-Op}}$. The following lemma shows that admissible strategies do not guarantee value-preserving property.
@@
Unlike the qualitative setting (Faella 2009; Aminof et al. 2020), Lemma 1 shows that we cannot characterize admissible strategies solely on the basis of $\operatorname{\mathbf{WCO}}$ strategies as they are not value-preserving in our quantitative setting. In the next section, we discuss two new categories of strategies derived from $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ that are always admissible. We then discuss their properties and show that they are also necessary conditions for admissibility. Table 1 summarizes properties of all new strategies we define hereafter.
@@
For every history $h$, we observe that strategies that are cooperative optimal have the least payoff. Thus, every $\sigma$ that is $\operatorname{\mathbf{Co-Op}}$ is admissible as there does not exist $\sigma^{\prime}$ that weakly dominates it. We now define strongly cooperative condition ($\operatorname{\mathbf{SC}}$) which generalizes $\operatorname{\mathbf{Co-Op}}$. Intuitively, strategies that are $\operatorname{\mathbf{SC}}$ have a lower payoff than the worst-case optimal payoff at $h$. In case, a lower payoff cannot be obtained, $\operatorname{\mathbf{SC}}$ are worst-case optimal.
@@
winning
Value-preserving
Winning
Memoryless
Algorithm Sec. 3 Sec. 3 Sec. 4 Sec. 3 Sec. 4
Table 1: Properties of strategies
From Lemma 2, it suffices for us to show that $\operatorname{\mathbf{SC}}$ strategies always exist to prove that admissible strategies always exist. Unfortunately, $\operatorname{\mathbf{SC}}$ strategies are history-dependent, i.e., we need to reason over every state along a history to check for admissibility.
@@
We now look at another interesting class of strategy that is always admissible. For every history $h$, $\operatorname{\mathbf{WCO}}$ strategies always guarantee the worst-case payoff. A subset of $\operatorname{\mathbf{WCO}}$ strategies that are also cooperative optimal ensures that such strategies enforce a worst-case payoff while safely minimizing the payoff. We call such strategies Worst-case Cooperative Optimal strategies ($\operatorname{\mathbf{WCo-Op}}$).
@@
where $\operatorname{acVal}(h):=\min\{\operatorname{cVal}(h,\sigma^{\prime})|\sigma^{\prime}\in\Sigma,\operatorname{aVal}(h,\sigma)\leq\operatorname{aVal}(h)\}$ is the optimal adversarial-cooperative value of $h$.
In the game in Fig. 2, only $\sigma_{2}$ is $\operatorname{\mathbf{WCo-Op}}$ strategy. That is because action $v_{4}\to v_{7}$ belongs to an admissible strategy as it has the minimum cooperative value while ensuring the worst-case optimal payoff. Here $\operatorname{aVal}(v_{4},\sigma_{2})=\operatorname{aVal}(v_{4})=9$; $\operatorname{cVal}(v_{4},\sigma_{2})=2$.
@@
From Lemma 3, it suffices for us to show that $\operatorname{\mathbf{WCo-Op}}$ strategies always exist for admissible strategies to always exist. Interestingly, unlike prior work (Brenguier et al. 2016), in our case, $\operatorname{\mathbf{WCo-Op}}$ strategies do always exist and at least one is memoryless. The latter is desirable because memoryless strategies can be computed efficiently using fixed-point like algorithms (Baier and Katoen 2008).
@@
A consequence of Thm. 2 is that a subset of admissible strategies, precisely $\operatorname{\mathbf{WCo-Op}}$ strategies, are history-independent even for a payoff that is history-dependent. Thus, for $\sigma$ to be admissible, it is sufficient to be $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$, i.e.,
@@
As shown below (Thm. 7), $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ are also necessary conditions for admissibility. This becomes useful for synthesizing the set of all admissible strategies.
@@
Analogously, a strategy $\sigma$ is not admissible if there exists another strategy $\sigma^{\prime}$ that weakly dominates it. Formally, $\sigma^{\prime}\succ\sigma$ if, and only if, $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$, the following holds:
Lemma 4.
@@
There always exists an admissible strategy in a 2-player, turn-based, total-payoff, reachability game $\mathcal{G}$.
Observe that Thm. 7 characterizes the set of all admissible strategies and Thm. 4 establishes their existence in full generality for 2-player turn-based games with reachability objectives. For computability consideration, we bound the payoffs associated with plays to a given budget $\mathcal{B}$ so that the set of all admissible strategies is finite. This is a reasonable assumption as, in many real-world applications, systems usually have limited resources like battery.
@@
Given $\mathcal{G}$ and budget $\mathcal{B}$, we first construct a game tree arena that captures all plays with payoff less than or equal to $\mathcal{B}$. Next, we show that the payoff function on this tree is history-independent, which allows us to modify Thm. 7 and compute a finite set of $\operatorname{aVal}$s. We first conclude the following from Lemma 2 and Thm 1.
@@ 11                   $\Sigma_{adm}:h\to v^{\prime}$ Add only states in $h$
Algorithm 1 Pseudocode for Problem 1
@@
A consequence of Corollary 1 is that we cannot use a backward induction-based algorithm to compute these strategies. Thus, we use a forward search algorithm that starts from the initial state and recursively checks whether the admissibility constraints are satisfied.
@@
The algorithm is outlined in Alg. 1. Given game $\mathcal{G}$ and budget $\mathcal{B}$, we construct a tree of plays $\mathcal{G}^{\prime}$ by unrolling $\mathcal{G}$ until the payoff associated with a play exceeds $\mathcal{B}$ or a goal state is reached. Every play in $\mathcal{G}$ corresponds to a branch in $\mathcal{G}^{\prime}$. Every play that reaches a goal state with a payoff $b\leq\mathcal{B}$ in $\mathcal{G}$ is a play that ends in a leaf node in $\mathcal{G}^{\prime}$, which is marked as a goal state for Sys player. The leaf nodes in $\mathcal{G}^{\prime}$ that correspond to the plays with payoff $b>\mathcal{B}$ in $\mathcal{G}$ are assigned a payoff of +$\infty$+ ∞. Further, in $\mathcal{G}^{\prime}$, the weights along all the edges are zero, and the payoffs are strictly positive only when a play reaches a leaf node, otherwise it is zero. By construction, the payoff function $\operatorname{Val}$ is history-independent in $\mathcal{G}^{\prime}$. Informally, a history-independent $\operatorname{Val}$ assigns values to plays according to only the current state and does not depend on the history of visited states. Then, Thm. 7 can be restated as the following lemma.
@@
Given $\mathcal{G}^{\prime}$, strategy $\sigma$ is admissible if, and only if, $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$ with $\operatorname{last}(h)\in V_{s}$ and $v^{\prime}=\delta(\operatorname{last}(h),\sigma(h))$, the following holds,
@@
$\big{(}\operatorname{aVal}(\operatorname{last}(h))=\operatorname{aVal}(v^{\prime})=\operatorname{cVal}(v^{\prime})=\operatorname{acVal}(\operatorname{last}(h))\big{)}$ (9b)
where $\operatorname{aValues}:=\{\operatorname{aVal}(v)|v\in h\}$ is the set of adversarial values along history $h$.
In Alg. 1, we first construct $\mathcal{G}^{\prime}$ as discussed before. Then, we compute $\operatorname{aVal}$ and $\operatorname{cVal}$ values for each state in $\mathcal{G}^{\prime}$ using the Value Iteration algorithm (Brihaye et al. 2017). If $\mathcal{B}<\operatorname{cVal}(v_{0})$, then there does not exist a play that reaches a goal state in $\mathcal{G}^{\prime}$. Thus, all strategies in $\mathcal{G}$ are admissible. To compute admissible strategies on $\mathcal{G}^{\prime}$, we use a DFS algorithm to traverse every play and check if the admissibility criteria from Lemma 5 is satisfied. If yes, we add the history and the successor state to $\Sigma_{adm}$. We repeat this until every state in $\mathcal{G}^{\prime}$ is explored. This algorithm is sound and complete with polynomial time complexity.
@@
Given $\mathcal{G}$ and $\mathcal{B}$, Algorithm 1 returns the set of all admissible strategies $\Sigma_{adm}$ and has polynomial time complexity.
@@ 4 Admissible Winning Strategies
In Sec. 3, we show that $\operatorname{\mathbf{SC}}$ strategies are willing to risk a higher payoff $(\operatorname{aVal}(\sigma(h))>\operatorname{cVal}(h,\sigma))$ in the hopes that the Env will cooperate, i.e., they are optimistic.
@@
Lemma 7 (Existence of admissible winning strategies).
There always exists an admissible winning strategy in a 2-player, turn-based, total-payoff, reachability games $\mathcal{G}$.
Figure 3: Top Row: (b) $\sigma_{adm}$ for $\mathcal{B}=5$ and (c) $\sigma^{win}_{adm}$ for $\mathcal{B}=12$. Bottom Row $\mathcal{B}=12$: (e) $\sigma_{adm}$ and (f) $\sigma^{win}_{adm}$.
@@
Here, we give the algorithm to solve Problem 2. Similar to admissible strategies, admissible winning strategies are history-dependent as shown by the next theorem.
@@
Thus, memoryless strategies are not sufficient for strategies to be admissible winning. The algorithm is the same as Alg. 1 except for checking the admissibility criteria. The construction of the tree, the syntax, and semantics are as outlined in Sec. 3. We modify the admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ This modification expresses the value-preserving term from Eq. (10a).
@@
Given $\mathcal{G}$ and $\mathcal{B}$, the algorithm described returns the set of all admissible winning strategies $\Sigma^{win}_{adm}$.
5 Examples

### v2

@@ 3 Admissible Strategies
Note $V_{win}$, $V_{los}$, and $V_{pen}$ define a partition for $V$, i.e., their union is $V$ and their pair-wise intersection is the empty set. Based on these sets, we characterize value-preserving strategies according to the notion of value for each state. Let $\operatorname{sVal}:V\to\{-1,0,1\}$ be a state-value function such that $\operatorname{sVal}(v)=1$ if $v\in V_{win}$, 00 if $v\in V_{pen}$, and $-1$ if $v\in V_{los}$.
@@
We say history $h$ is value-preserving if $\operatorname{sVal}(h_{j})\leq\operatorname{sVal}(h_{j+1})$ for all $10\leq j<|h|-$ 1. Strategy $\sigma$ is value preserving if every $h\in\operatorname{Plays}^{v_{0}}(\sigma)$ is value preserving.
Let us now look at two classical notions for strategies defined for quantitative games and discuss their value-preserving property. We say $\sigma$ is a worst-case optimal strategy ($\operatorname{\mathbf{WCO}}$) if $\operatorname{aVal}(h,\sigma)=\operatorname{aVal}(h)$. If at the current state $v$, $\operatorname{sVal}(v)=1$, then an optimal winning strategy $\sigma_{win}$ exists such that all plays in $\operatorname{Plays}^{v}(\sigma_{win})$ are value-preserving. Since $\sigma_{win}$ is $\operatorname{\mathbf{WCO}}$, every $\operatorname{\mathbf{WCO}}$ strategy in $V_{win}$ is also value preserving. If $\operatorname{sVal}(v)\neq-1$, a cooperatively-optimal ($\operatorname{\mathbf{Co-Op}}$) strategy $\sigma$ exists such that $\operatorname{cVal}(h,\sigma)=\operatorname{cVal}(h)$. Unlike $\operatorname{\mathbf{WCO}}$, $\operatorname{\mathbf{Co-Op}}$ strategies are not value-preserving. In Fig. 2, $\sigma_{1}$ and $\sigma_{2}$ are $\operatorname{\mathbf{WCO}}$ as they ensure the lowest payoff of 10 if Env is adversarial while $\sigma_{3}$ is $\operatorname{\mathbf{Co-Op}}$ as the corresponding payoff of 1 is the lowest for a cooperative Env. Further, notice that $\sigma_{3}$, while admissible, is not value preserving. The following lemma formalizes this observation.
@@
Unlike the qualitative setting [Faella, 2009; Aminof et al., 2020], Lemma 1 shows that we cannot characterize admissible strategies solely on the basis of $\operatorname{\mathbf{WCO}}$ strategies as they are not value-preserving in our quantitative setting. Below, we derive two new categories of strategies from $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ that are always admissible. We then discuss their properties and show that they are also necessary conditions for admissibility. Table 1 summarizes properties of all new strategies we define hereafter.
winning
Value-preserving
Winning
Memoryless
Algorithm Sec. 3 Sec. 3 Sec. 4 Sec. 3 Sec. 4
Table 1: Properties of new strategies defined for admissibility.
@@
Note that, for every history $h$, the strategies that are cooperative optimal have the least payoff. Thus, every $\sigma$ that is $\operatorname{\mathbf{Co-Op}}$ is admissible as there does not exist $\sigma^{\prime}$ that weakly dominates it. We now define strongly cooperative condition ($\operatorname{\mathbf{SC}}$) which generalizes $\operatorname{\mathbf{Co-Op}}$. Intuitively, strategies that are $\operatorname{\mathbf{SC}}$ have a lower payoff than the worst-case optimal payoff at $h$. In case, a lower payoff cannot be obtained, $\operatorname{\mathbf{SC}}$ are worst-case optimal.
@@
From Lemma 2, it suffices for us to show that $\operatorname{\mathbf{SC}}$ strategies always exist to prove that admissible strategies always exist. Unfortunately, $\operatorname{\mathbf{SC}}$ strategies are history-dependent, i.e., we need to reason over every state along a history to check for admissibility. This is formalized as follows.
@@
We now look at another interesting class of strategy that is always admissible. For every history $h$, $\operatorname{\mathbf{WCO}}$ strategies always guarantee the worst-case payoff. A notable subset of $\operatorname{\mathbf{WCO}}$ strategies are those that are also $\operatorname{\mathbf{Co-Op}}$. We call such strategies Worst-case Cooperative Optimal ($\operatorname{\mathbf{WCo-Op}}$).
@@
where $\operatorname{acVal}(h):=\min\{\operatorname{cVal}(h,\sigma)\mid\sigma\in\Sigma,\operatorname{aVal}(h,\sigma)\leq\operatorname{aVal}(h)\}$ is the optimal adversarial-cooperative value of $h$.
In this definition, $\operatorname{acVal}$ is a new notion that characterizes the minimum payoff that Sys player can obtain from the set of worst-case optimal strategies. In the game in Fig. 2, only $\sigma_{2}$ is $\operatorname{\mathbf{WCo-Op}}$ strategy. That is because action $v_{4}\to v_{7}$ belongs to an admissible strategy as it has the minimum cooperative value while ensuring the worst-case optimal payoff. Here $\operatorname{aVal}(v_{4},\sigma_{2})=\operatorname{aVal}(v_{4})=9$; $\operatorname{cVal}(v_{4},\sigma_{2})=2$.
@@
From Lemma 3, it suffices for us to show that $\operatorname{\mathbf{WCo-Op}}$ strategies always exist for admissible strategies to always exist. Interestingly, unlike prior work [Brenguier et al., 2016], in our case, $\operatorname{\mathbf{WCo-Op}}$ strategies always exist and at least one is memoryless. The latter is desirable because memoryless strategies can be computed efficiently using fixed-point-based algorithms [Baier and Katoen, 2008].
@@
The proof relies on the fact that memoryless strategies are sufficient for $\operatorname{aVal}$ and $\operatorname{cVal}$, and proceeds by constructing a subgame $\bar{\mathcal{G}}$ for a given history $h$ and showing that the cooperative value of the initial state in $\bar{\mathcal{G}}$ equals the $\operatorname{acVal}(h)$ of $\mathcal{G}$. A consequence of Theorem 2 is that a subset of admissible strategies, precisely $\operatorname{\mathbf{WCo-Op}}$ strategies, are history-independent even for a payoff that is history-dependent. Thus, for $\sigma$ to be admissible, it is sufficient to be $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$, i.e.,
@@
As shown below (Theorem 3), $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ are also necessary conditions for admissibility. This becomes useful for synthesizing the set of all admissible strategies.
@@
[…]
@@
There always exists an admissible strategy $\sigma_{adm}$ in a 2-player, turn-based, total-payoff, reachability game $\mathcal{G}$, and $\sigma_{adm}$ solves Problem 1 if $\operatorname{cVal}(v_{0},\sigma_{adm})\leq\mathcal{B}$.
The proof follows from Lemma 3 and Theorem 2. Observe that Theorem 3 characterizes the set of all admissible strategies and Theorem 4 establishes their existence in full generality, i.e., independent of $\mathcal{B}$, for 2-player turn-based games with reachability objectives. For computability considerations, we bound the payoffs associated with plays to a given budget $\mathcal{B}$ so that the set of all admissible strategies is finite. This is a reasonable assumption that is often used in, e.g., energy games with fixed initial credit and robotics application with finite resources [Chakrabarti et al., 2003; Bouyer et al., 2008; He et al., 2017; Filiot et al., 2010].
@@
Given $\mathcal{G}$ and budget $\mathcal{B}$, we first construct a game tree arena that captures all plays with payoff less than or equal to $\mathcal{B}$. Next, we show that the payoff function on this tree is history-independent, which allows us to modify Theorem 3 and compute a finite set of $\operatorname{aVal}$s. We conclude the following from Lemma 2 and Theorem 1.
@@ 11                   $\Sigma_{adm}:h\to v^{\prime}$ Add only states in $h$
Algorithm 1 Admissible Strategy Synthesis
@@
A consequence of Corollary 1 is that we cannot use a backward induction-based algorithm to compute these strategies. Instead, we use a forward search algorithm that starts from the initial state and recursively checks whether the admissibility constraints are satisfied.
@@
The algorithm is outlined in Alg. 1. Given game $\mathcal{G}$ and budget $\mathcal{B}$, we construct a tree of plays $\mathcal{G}^{\prime}$ by unrolling $\mathcal{G}$ until the payoff associated with a play exceeds $\mathcal{B}$ or a goal state is reached. Every play in $\mathcal{G}$ corresponds to a branch in $\mathcal{G}^{\prime}$. Every play that reaches a goal state with a payoff $b\leq\mathcal{B}$ in $\mathcal{G}$ is a play that ends in a leaf node in $\mathcal{G}^{\prime}$, which is marked as a goal state for Sys player. The leaf nodes in $\mathcal{G}^{\prime}$ that correspond to the plays with payoff $b>\mathcal{B}$ in $\mathcal{G}$ are assigned a payoff of +$\infty$+ ∞. Further, in $\mathcal{G}^{\prime}$, the weights along all the edges are zero, and the payoffs are strictly positive only when a play reaches a leaf node, otherwise it is zero. By construction, the payoff function $\operatorname{Val}$ is history-independent in $\mathcal{G}^{\prime}$. Then, Theorem 3 can be restated as the following lemma.
@@
Given $\mathcal{G}^{\prime}$, strategy $\sigma$ is admissible if and only if $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$ with $\operatorname{last}(h)\in V_{s}$ and $v^{\prime}=\delta(\operatorname{last}(h),\sigma(h))$, the following holds,
@@
$\big{(}\operatorname{aVal}(v^{{\dagger}})=\operatorname{aVal}(v^{\prime})=\operatorname{cVal}(v^{\prime})=\operatorname{acVal}(v^{{\dagger}})\big{)}$ (8b)
where $\operatorname{aValues}:=\{\operatorname{aVal}(v)\mid v\in h\}$ is the set of adversarial values along history $h$ and $v^{{\dagger}}=\operatorname{last}(h)$.
After constructing $\mathcal{G}^{\prime}$, Alg. 1 computes $\operatorname{aVal}$ and $\operatorname{cVal}$ values for each state in $\mathcal{G}^{\prime}$ using the Value Iteration algorithm [Brihaye et al., 2017]. If $\mathcal{B}<\operatorname{cVal}(v_{0})$, then there does not exist a play that reaches a goal state in $\mathcal{G}^{\prime}$. Thus, all strategies in $\mathcal{G}$ are admissible. To compute admissible strategies on $\mathcal{G}^{\prime}$, we use a DFS algorithm to traverse every play and check if the admissibility criteria from Lemma 4 is satisfied. If yes, we add the history and the successor state to $\Sigma_{adm}$. We repeat this until every state in $\mathcal{G}^{\prime}$ is explored. This algorithm is sound and complete with polynomial time complexity.
@@
Given $\mathcal{G}$ and a fixed budget $\mathcal{B}$, Alg. 1 returns the set of all admissible strategies $\Sigma_{adm}$ and has polynomial time complexity.
@@ 4 Admissible Winning Strategies
In Sec. 3, we show that $\operatorname{\mathbf{SC}}$ strategies are willing to risk a higher payoff $(\operatorname{aVal}(\delta(\operatorname{last}(h),\sigma(h)))>\operatorname{cVal}(h,\sigma))$ in the hopes that the Env will cooperate, i.e., they are optimistic.
@@
Lemma 6.
There always exists an admissible winning $\sigma^{win}_{adm}$ strategy in a 2-player, turn-based, total-payoff, reachability games $\mathcal{G}$, and $\sigma^{win}_{adm}$ solves Problem 2 if $\operatorname{aVal}(v_{0},\sigma^{win}_{adm})\leq\mathcal{B}$.
@@
Here, we give an algorithm to solve Problem 2. Similar to admissible strategies, admissible winning strategies are history-dependent as shown by the next theorem.
@@
The algorithm is similar to Alg. 1 except for the admissibility checking criteria. The tree construction, syntax, and semantics are as outlined in Sec. 3. We modify admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ This modification expresses the value-preserving term from Eq. (9a).
@@
Given $\mathcal{G}$ and a fixed budget $\mathcal{B}$, the algorithm described returns the set of all admissible winning strategies $\Sigma^{win}_{adm}$.
5 Illustrative Examples